the longest side of a right triangle is 7 ft longer than four times the length of the shortest side. the other side of the triangle is 7 ft shorter than four times the length of the shortest side. find the lengths of the sides of the triangle.
Answer by amarjeeth123(569)

Let the shortest side be x.
The longest side is (7+4x).
The other side is (4x-7).
We have (4x+7)^2=(4x-7)^2+x^2
16x^2+56x+49=16x^2-56x+49+x^2
x^2=112x
x=112.
The shortest side is 112.
The longest side is 455.
The other side is 441.